OpenStudy (anonymous):

What is 5 pi over 3 radians in degrees?

OpenStudy (***[isuru]***):

Hi, \[\pi = 180^{0}\] so \[\frac{ 5\pi }{ 3 } = \frac{ 5\times180 }{ 3 } = 300^{0}\] hope this will help ya !!!!